OpenStudy (anonymous):

How do you find the solution to this problem: if you have 90 green beads, and 108 blue beads what is the greatest # of identical necklaces you can make using all the beads?

OpenStudy (anonymous):

is there any more information to the question? if not then the only way you could make the maximum number of necklaces is to have 1 green bead and 1 blue bead on each necklace, thus the maxium possible amount of necklaces will be when you run out of green beads, so 90 would be my answer

OpenStudy (anonymous):

I believe it is a LCM problem. When I looked online for answer I got 18 -however, I am not how to reach that solution.

OpenStudy (anonymous):

if its lcm, then i'd write it out like this: lcm(90,108) so focussing on the 90 i would do... 90/2 = 45 store 2 for now 45 is 9*5, so now we have 2 and 5 and 9 is 3*3, so now we have 2*5*3*3 = 90 or 2*3^2*5 = 90 then doing the same for 108 i get 108/2 = 54, 54/2 = 27, 27 = 3*9, 9=3*3, then 2*2*3*9 =108 which is 2*2*3*3*3 which is 2^2 * 3^3 now combining both of them 2*3^2*5 2^2 * 3^3

OpenStudy (anonymous):

2*3^2*5 = 90 2^2 * 3^3=108 so we have 2^2 * 3^3 * 5 = 540